MICROWAVE SPECTRUM AND PARTIAL GAS PHASE STRUCTURE OF A FORMIC ACID-FORMAMIDE COMPLEX ^{a}

\underline{ADAM M. DALY}, STEPHEN G. KUKOLICH, Department of Chemistry and Biochemistry, 1306 E. University, Tucson, Arizona 85721; BRYAN SARGUS, Catalina Foothills High School Teacher, Tucson, Arizona 85718.

The rotational spectrum of 5 isotopomers of the complex formed between formic acid and formamide have been measured and analyzed using a pulsed beam fourier transform microwave spectrometer. The rotational constants and nitrogen’s quadrupole coupling strength for HC^{12}OOH − H_{2}N^{14}COH are A=5889.51(9), B=2148.734(2), 1575.127(1), eQq_{aa} = 1.020(8), eQq_{bb} = 1.98(1) and eQq_{cc} = −2.998(9) MHz. Using the 15 rotational constants obtained with HC^{13}OOH, HCOOD, DCOOH and H_{2}N^{15}CHO isotopologes, key structural information was obtained from a least squares fit. R(CO-HN) distance of 1.77Å and R(OH-OC) distance of 1.81Å and an (COH) angle of formic acid of 121^{∘} was obtained that was very different from the monomer value of 106.9^{∘}. Density functional theory using B3PW91, HCTH407 and TPSS and MP2 calculations were performed using 6-311++G(d,p) and compared to experimentally determined values.

aSupported by THE NATIONAL SCIENCE FOUNDATION